OpenStudy (anonymous):

Use the following problem to answer Max P = 65x+25y 10x + 16y ≥6000 10x +8y ≤ 3600 x,y ≥

OpenStudy (anonymous):

So far I have 10x +16y = 6000 From (0,0) I got points (0,375), (600,0) 10x+8y = 3600 From (0,0) I got pints (0,450) (360,0) So 10( 0) +16(0) ≥ 6000 - NO 10(0) +8(0) ≤ 36000 YES NEED HELP!!! I have 10x +16y ≥ 6000 I multiplied 10x+8y ≤ 3600 by 2 giving me -20x - 16y ≤ - 7200 x= 720 BUT - none of my answers start with 720 SO what am I doing wrong??!

OpenStudy (anonymous):

wow I was doing the math wrong! I actually get x = 120!!! and I plug 120 into my second equation - is that right anyone?